OpenStudy (anonymous):

Select the phrase that correctly completes the sentence. A donut box (on the left) and a cereal box (on the right) are placed side-by-side. The donut box has length thirteen inches, width nine inches, and height three inches. The cereal box has length ten inches, width two inches, and height fourteen inches. The volume of the donuts box (on the left) is blank the volume of the cereal box (on the right). the same as 3 cubic inches less than 6 cubic inches less than seventy-one cubic inches more than don't understand please help

OpenStudy (anonymous):

71 cubic inches more than.

OpenStudy (anonymous):

13*9*3 = 351 10*2*14 = 280 351 - 280 = 71

OpenStudy (anonymous):

thanks